# Simulating Roman+Gaia Combined Astrometry, Parallaxes, and Proper Motions

### v1

@@ I Introduction
Fortunately, the next generation of astrometry – largely based on active data collection from Euclid (Euclid Collaboration et al., 2025) and Rubin/LSST, (Ivezić et al., 2019) and the soon-to-launch Roman Space Telescope (Spergel et al., 2015; Akeson et al., 2019) – will complement Gaia’s successes by extending to much fainter magnitudes with similar precision in proper motion (PM) and parallax. While high precision stellar astrometry is not the primary goal of many of these missions/telescopes, they will nonetheless deliver incredible stellar localizations down to very faint sources (e.g., mag for source detections for many Roman surveys; Observations Time Allocation Committee and Community Survey Definition Committees, 2025). These new faint stars will increase the number of stars in a given population by more than an order of magnitude, unlocking a wealth of previously hidden information. A far-from-exhaustive list of potential science applications includes: high spatial and distance resolution views of MW dynamics to understand its merger history (e.g., Bullock and Johnston, 2005; Antoja et al., 2018; Cunningham et al., 2019; Apfel et al., 2025), the discovery and characterization of distant MW stellar halo kinematic substructure (e.g., Belokurov et al., 2018; Helmi et al., 2018; Naidu et al., 2020), detection of faint or dark companions in low-inclination binary systems (e.g., Gaia Collaboration et al., 2024), bulk motions of nearby galaxies to understand the Local Group interaction history (e.g., van der Marel et al., 2012; McConnachie and Venn, 2020; Pace et al., 2022; Bennet et al., 2024), and resolved internal kinematics of MW stellar streams (e.g., Ibata et al., 2020; Bonaca and Price-Whelan, 2025; Nibauer et al., 2025) and nearby dwarf galaxies (e.g., Read et al., 2021; Massari et al., 2018; Vitral et al., 2024, 2025) for constraints on the nature of dark matter. In this paper, we specifically focus on the capabilities of Roman, so we point to the work of WFIRST Astrometry Working Group et al. (2019), Han et al. (2023), and Dey et al. (2023) for a more-complete list of the Local Group science applications that pristine Roman astrometry will significantly advance. It is difficult to understate the inevitable scientific breakthroughs that will be enabled by next generation high precision astrometry over the next 5 to 10 years.
@@
In preparing for the astrometric future, it is prudent to ensure that we harness the full potential of previous datasets and surveys, namely those of Gaia. While it is often easier to measure stellar PMs from images taken with the same instrument/filter/telescope orientation because it removes many sources of difficult-to-calibrate systematics (e.g. Anderson and King, 2004, 2006; Bellini and Bedin, 2009; Bellini et al., 2011; Anderson, 2007, 2022; Sohn et al., 2012, 2018), recent work has shown success with cross-observatory astrometry measurements. Various Gaia data releases, for example, have been combined with different telescopes and surveys – including Pan-STARRS1 (Magnier et al., 2020), HST (e.g., Massari et al., 2017; del Pino et al., 2022; Warfield et al., 2023; McKinnon et al., 2024), JWST (Libralato et al., 2023, 2024b), and Euclid (Libralato et al., 2024a) – to measure significantly more precise PMs by benefiting from an increased time baseline.
@@ II Underlying Theory & Statistics
We are interested in making 5 astrometric measurements per star: its true position on the sky at some reference epoch (2D measurement), its parallax (1D measurements), and its proper motion vector (2D measurement). We will make this measurement by comparing Gaia data – providing prior information about positions, parallaxes, and PMs – with positions measured in different Roman images at different epochs. Of course, almost all of the math presented here also applies to other telescopes, such as HST, JWST, Rubin, and Euclid. The statistics presented in this section are inspired by the math presented in the work describing the HST+Gaia astrometry tool BP3M (Bayesian Positions, Parallaxes, and Proper Motions; McKinnon et al., 2024).
We start by defining some key terms: we set to be the index of the star, to be the time of the reference epoch (i.e. J2016.0 for Gaia DR2 and DR3), and to be the true position of the star at time . Here, the uppercase refers to “True” while the lowercase refers to “time”. We note that is the parallax-independent position of the star, meaning it is the position of the star if it had zero parallax. Assuming the star’s distance from the Sun is not changing much over time – that is, the radial velocity of the star does not significantly change the heliocentric distance over the course of a few decades111As estimated in Appendix A, we require pc for a 100 km/s line-of-sight velocity over a 10 year time baseline to ensure the PM impact from changing distance is smaller than 0.01 mas/yr. – we have the following time evolution definition to describe the true position at time :
@@
where the subscript refers to Gaia (which is defined at reference epoch ), is the Gaia-measured parallax-independent position222One could argue that , , and in and should refer to the true parallax-independent position instead of the Gaia-measured values, but this would prevent the problem from having a clean linear solution. In any case, the difference is unlikely to matter significantly given the small Gaia uncertainties in position (i.e. mas). Given and , we can define while also measuring the PM and parallax for a star., and we have now chosen to set . The and definitions must contain terms that account for updates to the position measured by Gaia; that is, we expect that the Gaia position is slightly different from the truth, and we would like to update it to a more accurate value when all our measurements are combined. We note that the term could also be adjusted to include a constant acceleration vector (e.g., terms in ).
@@ III Astrometry Simulation Tool
The first step is to estimate the Gaia astrometric uncertainty, which is provided by the Gaia collaboration forecasting different precisions as a function of magnitude for DR3 through DR5 444Gaia astrometric precision relationships originally found at this webpage: https://www.cosmos.esa.int/web/gaia/science-performance. For posterity, we state those relationships here:
@@
Next, we use the pandeia555see pandeia documentation at https://roman-docs.stsci.edu/simulation-tools-handbook-home/roman-wfi-exposure-time-calculator/pandeia-for-roman/overview-of-pandeia (Pontoppidan et al., 2016) and stpsf666see stpsf documentation at https://stpsf.readthedocs.io/en/latest/intro.html (Perrin et al., 2012, 2025) tools to simulate how well we can expect to measure the position of a source in a Roman image as a function of filter, magnitude, and exposure time. We start by obtaining a super-sampled version of the Roman PSF for a given filter using stpsf. We then generate a simulated Roman observation of a star at a given flux level, including Poisson noise. Using the super-sampled PSF, we iteratively fit for the position of the star as well as its flux. In particular, we take numerical derivatives of the PSF with respect to the position to linearize the fitting problem, which allows us to quickly update our parameter guesses and converge on the final answer. These numerical derivatives, when combined with arguments about the shape of the fitting surface, also allow us to estimate the full covariance matrix that relates the position and flux uncertainties. Agreeing with the expectation from the Cramér–Rao bound (Cramér, 1946; Rao, 1947), we have:
@@
An example of these outputs are shown in Figure 2 for a second exposure. Increasing the exposure time typically has the effect of sliding these curves to the right, such that position uncertainties for fainter stars increase compared to shorter exposures, up to the limit for saturated sources. To be clear, the position uncertainty used in the final astrometry precision calculation is defined by profiles similar to the ones in Figure 2, which are then added in quadrature with an adjustable floor position uncertainty, nominally set to 1% of the Roman pixel width. From this figure, we see that F146 gives the most precise position per exposure for most magnitudes, likely because this wide filter is able to collect more flux SNR per exposure than its narrower counterparts. However, we caution that the large wavelength width of F146 may allow difficult-to-handle chromatic variations in the PSF to impact position measurements. While we may be able to quantify this effect once we start collecting real Roman data, chromatic PSF variation across F146 might ultimately lead to large systematic uncertainties that we do not account for in our simulations.
@@
Under the curves for the various Roman filters, we have drawn a thick blue line to show the expected trend in position uncertainty across different magnitude regimes, which is general to all telescopes/filters/exposure times. At the brightest end, pixels become saturated and cannot constrain better positions for more input flux; in fact, for many telescopes there is a magnitude limit where the position uncertainty is effectively infinite/undefined because localization becomes impossible when too many neighboring pixels saturate. Here, our saturation regime is defined by the pandeia flux SNR behavior. In the intermediate magnitudes, bright sources are dominated by Poisson noise from the source, meaning the following relationship is expected:
where the final line in the above equation comes from Equation 8. For faint sources, the background begins to dominate the noise and leads to faster increases in the position uncertainty, producing the following expected relationship:
@@ III.1 Example Application
We choose an arbitrary configuration as an illustrative example, setting the target coordinates to and , which is the approximate center of the HLWAS Medium Field 1. We choose to observe at three Roman epochs: J2028.0, J2031.0, and J2031.5. The first and final epochs are both set to use the F087 filter with 4 dithers per epoch, while the middle epoch is set to use the F106 filter with 5 dithers. For all epochs, each individual observation is set to use a second exposure defined by the multi-accumulation (MA) strategy IM_193_11777Please refer to the following webpage for a detailed description of the allowed Roman multi-accumulation stategies: https://roman-docs.stsci.edu/roman-instruments/the-wide-field-instrument/observing-with-the-wfi/wfi-multiaccum-ma-tables. We note that this MA is the one used to produce the position uncertainty functions shown in Figure 2. Finally, we assume that all stars have zero color in all pairwise filter combinations, including between the Roman and Gaia filters, and simulate stars with magnitudes in the range of mag. Users interested in a particular type of star (e.g., KIII giants or MSTO stars) could use stellar isochrones to determine more accurate colors and magnitudes in each filter.
@@ IV.1 Position Uncertainties
Our calculations assume there will not be significant uncertainties associated with aligning Roman images onto the Gaia reference frame, either through systematics or by simple propagation that leads to inflated localization uncertainties. That is, we assume the alignment uncertainty is a component of the floor uncertainty in our positions. Because of the large Roman field of view, this is likely a very reasonable choice: most pointings will contain enough well-measured Gaia stars (e.g., ) to align each image. However, there may be sparse regions – for example, in the HLWAS – where there are not many Gaia stars per Roman pointing. When properly propagated, this alignment uncertainty may increase the uncertainty of the Roman positions above what we have assumed.

### v2

@@ I Introduction
Fortunately, the next generation of astrometry – largely based on active data collection from Euclid (Euclid Collaboration et al., 2025) and Rubin/LSST, (Ivezić et al., 2019) and the soon-to-launch Roman Space Telescope (Spergel et al., 2015; Akeson et al., 2019) – will complement Gaia’s successes by extending to much fainter magnitudes with similar precision in proper motion (PM) and parallax. While high precision stellar astrometry is not the primary goal of many of these missions/telescopes, they will nonetheless deliver incredible stellar localizations down to very faint sources (e.g., mag for source detections for many Roman surveys; Observations Time Allocation Committee and Community Survey Definition Committees, 2025). These new faint stars will increase the number of stars in a given population by more than an order of magnitude, unlocking a wealth of previously hidden information. A far-from-exhaustive list of potential science applications includes: high spatial and distance resolution views of MW dynamics to understand its merger history (e.g., Bullock and Johnston, 2005; Antoja et al., 2018; Cunningham et al., 2019; Apfel et al., 2025), the discovery and characterization of distant MW stellar halo kinematic substructure (e.g., Belokurov et al., 2018; Helmi et al., 2018; Naidu et al., 2020), detection of faint or dark companions in low-inclination binary systems (e.g., Gaia Collaboration et al., 2024), bulk motions of nearby galaxies to understand the Local Group interaction history (e.g., van der Marel et al., 2012; McConnachie and Venn, 2020; Pace et al., 2022; Bennet et al., 2024), and resolved internal kinematics of MW stellar streams (e.g., Ibata et al., 2020; Bonaca and Price-Whelan, 2025; Nibauer et al., 2025) and nearby dwarf galaxies (e.g., Read et al., 2021; Massari et al., 2018; Vitral et al., 2024, 2025) for constraints on the nature of dark matter. In this paper, we specifically focus on the capabilities of Roman, so we point to the work of WFIRST Astrometry Working Group et al. (2019), Han et al. (2023), and Dey et al. (2023) for a more-complete list of the Local Group science applications that pristine Roman astrometry will significantly advance. For a focused discussion on Euclid’s potential impact on Local Group astrometry, please refer to Bedin (2025). It is difficult to understate the inevitable scientific breakthroughs that will be enabled by next generation high precision astrometry over the next 5 to 10 years.
@@
In preparing for the astrometric future, it is prudent to ensure that we harness the full potential of previous datasets and surveys, namely those of Gaia. While it is often easier to measure stellar PMs from images taken with the same instrument/filter/telescope orientation because it removes many sources of difficult-to-calibrate systematics (e.g. Anderson and King, 2004, 2006; Bellini and Bedin, 2009; Bellini et al., 2011; Anderson, 2007, 2022; Sohn et al., 2012, 2018), recent work has shown success with cross-observatory astrometry measurements. Various Gaia data releases, for example, have been combined with different telescopes and surveys – including Pan-STARRS1 (Magnier et al., 2020), HST (e.g., Massari et al., 2017; del Pino et al., 2022; Warfield et al., 2023; McKinnon et al., 2024), JWST (Libralato et al., 2023, 2024b), and Euclid (Libralato et al., 2024a; Bedin, 2025) – to measure significantly more precise PMs by benefiting from an increased time baseline.
@@ II Underlying Theory & Statistics
We are interested in making 5 astrometric measurements per star: its true position on the sky at some reference epoch (2D measurement), its parallax (1D measurements), and its proper motion vector (2D measurement). We will make this measurement by comparing Gaia data – providing prior information about positions, parallaxes, and PMs – with positions measured in different Roman images at different epochs. The statistics presented in this section are inspired by the math presented in the work describing the HST+Gaia astrometry tool BP3M (Bayesian Positions, Parallaxes, and Proper Motions; McKinnon et al., 2024). While this paper focuses on Roman data in combination with the Gaia mission, the framework we present is generally applicable to combine (rectified and aligned111We note that the rectification and alignment of images to a global reference frame is often the most difficult step in high-precision astrometry, but we assume this has been done for us in the body of this paper.) positions from any current or future missions.
We start by defining some key terms: we set to be the index of the star, to be the time of the reference epoch (i.e. J2016.0 for Gaia DR2 and DR3), and to be the true position of the star at time . Here, the uppercase refers to “True” while the lowercase refers to “time”. The true position of the star consists of its right ascension, , and declination, . We note that is the parallax-independent position of the star, meaning it is the position of the star if it had zero parallax. Assuming the star’s distance from the Sun is not changing much over time – that is, the radial velocity of the star does not significantly change the heliocentric distance over the course of a few decades222As estimated in Appendix A, we require pc for a 100 km/s line-of-sight velocity over a 10 year time baseline to ensure the PM impact from changing distance is smaller than 0.01 mas/yr. – we have the following time evolution definition to describe the true position at time :
@@
where the subscript refers to Gaia (which is defined at reference epoch ), is the Gaia-measured parallax-independent position333One could argue that , , and in and should refer to the true parallax-independent position instead of the Gaia-measured values, but this would prevent the problem from having a clean linear solution. In any case, the difference is unlikely to matter significantly given the small Gaia uncertainties in position (i.e. mas). Given and , we can define while also measuring the PM and parallax for a star., and we have now chosen to set . The and definitions must contain terms that account for updates to the position measured by Gaia; that is, we expect that the Gaia position is slightly different from the truth, and we would like to update it to a more accurate value when all our measurements are combined. We note that the term could also be adjusted to include a constant acceleration vector (e.g., terms in ).
@@ III Astrometry Simulation Tool
The first step is to estimate the Gaia astrometric uncertainty, which is provided by the Gaia collaboration forecasting different precisions as a function of magnitude for DR3 through DR5 555Gaia astrometric precision relationships originally found at this webpage: https://www.cosmos.esa.int/web/gaia/science-performance. For posterity, we state those relationships here:
@@
Next, we use the pandeia666see pandeia documentation at https://roman-docs.stsci.edu/simulation-tools-handbook-home/roman-wfi-exposure-time-calculator/pandeia-for-roman/overview-of-pandeia (Pontoppidan et al., 2016) and stpsf777see stpsf documentation at https://stpsf.readthedocs.io/en/latest/intro.html (Perrin et al., 2012, 2025) tools to simulate how well we can expect to measure the position of a source in a Roman image as a function of filter, magnitude, and exposure time. We start by obtaining a super-sampled version of the Roman PSF for a given filter using stpsf. We then generate a simulated Roman observation of a star at a given flux level, including Poisson noise. Using the super-sampled PSF, we iteratively fit for the position of the star as well as its flux. In particular, we take numerical derivatives of the PSF with respect to the position to linearize the fitting problem, which allows us to quickly update our parameter guesses and converge on the final answer. These numerical derivatives, when combined with arguments about the shape of the fitting surface, also allow us to estimate the full covariance matrix that relates the position and flux uncertainties. Agreeing with the expectation from the Cramér–Rao bound (Cramér, 1946; Rao, 1947), we have:
@@
An example of these outputs are shown in Figure 2 for a second exposure. Increasing the exposure time typically has the effect of sliding these curves to the right, such that position uncertainties for fainter stars decrease compared to shorter exposures, up to the limit for saturated sources. To be clear, the position uncertainty used in the final astrometry precision calculation is defined by profiles similar to the ones in Figure 2, which are then added in quadrature with an adjustable floor position uncertainty, nominally set to 1% of the Roman pixel width. From this figure, we see that F146 gives the most precise position per exposure for most magnitudes, likely because this wide filter is able to collect more flux SNR per exposure than its narrower counterparts. However, we caution that the large wavelength width of F146 may allow difficult-to-handle chromatic variations in the PSF to impact position measurements. While we may be able to quantify this effect once we start collecting real Roman data, chromatic PSF variation across F146 might ultimately lead to large systematic uncertainties that we do not account for in our simulations.
@@
Under the curves for the various Roman filters, we have drawn a thick blue line to show the expected trend in position uncertainty across different magnitude regimes, which is general to all telescopes/filters/exposure times. At the brightest end, pixels become saturated and cannot constrain better positions for more input flux; in fact, for many telescopes there is a magnitude limit where the position uncertainty is effectively infinite/undefined because localization becomes impossible when too many neighboring pixels saturate. Here, our saturation regime is defined by the pandeia flux SNR behavior.
Roman detectors, like those of JWST , are different from classical CCDs in that they collect time series counts in each pixel. The flux in a pixel is then calculated using the best-fit slope of the observed counts, sometimes referred to as “fitting a ramp” (e.g., Fixsen et al., 2000; Offenberg et al., 2001; Rauscher et al., 2007; Brandt, 2024; Li et al., 2026). One key benefit of these detectors is that they enable higher dynamic range images because any pixel that does not saturate within the first two readouts (e.g., seconds for most Roman readout strategies) is able to measure a scientifically useful flux888In theory, the first readout alone can be used to measure the flux. In practice, the resetting/flushing of the detector between exposures is imperfect such that counts from the previous exposure can bias the flux estimate.. However, the particular flux levels will determine which reads are unsaturated and useful for ramp fitting, which causes the flux S/N and position uncertainty to change as a non-monotonic function of magnitude. This likely explains the oscillating/saw-tooth behavior of the pandeia S/N outputs at the bright end of Figure 2.
We can split Roman pixels into unsaturated, “soft” saturated, and “hard” saturated regimes. The unsaturated regime is where all readouts remain unsaturated and can be used for measuring a flux; this threshold is filter and readout-pattern dependent, but most Roman filters appear completely unsaturated for mag in our 193 second exposure example (i.e. at the transition from “Saturated Source” to “Bright Source”). We define hard saturation as the point when fewer than two readouts are unsaturated such that ramp fitting becomes biased or impossible; again, this threshold is filter dependent but is mag for most Roman filters and exposure times999Single pixel saturation times for different Roman filters can be found here: https://science.nasa.gov/mission/roman-space-telescope/time-to-saturation-for-point-source-imaging/. Soft saturation occurs at intermediate magnitude ranges where one or more of the later readouts are saturated, but there are enough unsaturated readouts to constrain a useful flux measurement. It is the soft saturation regime that is highlighted by the “Saturated Source” label in Figure 2, and we do not attempt to implement a hard saturation threshold. Stars that produce hard saturated pixels will likely need to mask their cores when PSF fitting such that only the lower S/N pixels in the PSF wings will constrain positions. For Roman filters with very under-sampled PSFs (i.e., the bluer filters), masking the central pixel(s) will lead to a dramatic increase in position uncertainty because the wings contain relatively small amounts of position information. We leave a detailed analysis of the relationship between position uncertainty and the effects of changing numbers of soft and hard saturated pixels as future work that will be useful to understand the number of well-measured Gaia stars in each Roman image.
In the unsaturated intermediate magnitudes – that is, the range in Figure 2– bright sources are dominated by Poisson noise from the source, meaning the following relationship is expected:
where the final line in the above equation comes from Equation 8. For faint sources ( mag in Figure 2), the background begins to dominate the noise and leads to faster increases in the position uncertainty, producing the following expected relationship:
@@ III.1 Example Application
We choose an arbitrary configuration as an illustrative example, setting the target coordinates to and , which is the approximate center of the HLWAS Medium Field 1. We choose to observe at three Roman epochs: J2028.0, J2031.0, and J2031.5. The first and final epochs are both set to use the F087 filter with 4 dithers per epoch, while the middle epoch is set to use the F106 filter with 5 dithers. For all epochs, each individual observation is set to use a second exposure defined by the multi-accumulation (MA) strategy IM_193_11101010Please refer to the following webpage for a detailed description of the allowed Roman multi-accumulation stategies: https://roman-docs.stsci.edu/roman-instruments/the-wide-field-instrument/observing-with-the-wfi/wfi-multiaccum-ma-tables. We note that this MA is the one used to produce the position uncertainty functions shown in Figure 2. Finally, we assume that all stars have zero color in all pairwise filter combinations, including between the Roman and Gaia filters, and simulate stars with magnitudes in the range of mag. Users interested in a particular type of star (e.g., KIII giants or MSTO stars) could use stellar isochrones to determine more accurate colors and magnitudes in each filter.
@@ IV.1 Position Uncertainties
The PSF derivatives in Equation 7 suggest that subpixel shifts of the PSF distribution will impact the position uncertainties. We explore this effect in Figure 7 for three different filters: F062, F158, and F213. The widths of these PSFs change from very under-sampled (F062) to more complete sampling (F158 and F213). In all cases, we notice that the 2D PSF – and the 1D projections in X and Y – are asymmetric. The right-most column summarizes the impact that subpixel offsets have on the position uncertainty measurements, quantified by changes in . Given our pixel grid choices, stpsf places the center of the PSF in the middle of the central pixel. This explains why the bluest, most under-sampled filter shows a relatively large change from the zero-offset to (0.5,0.5) pixel offsets: the former has the PSF core falling in a single pixel while the latter splits the core light across fours pixels, enabling better centroiding. In this work, we report results using the zero-offset value (i.e. the bottom left corner of each right panel), as this is the default stpsf output. Future versions of our tool will likely let users select values for different subpixel offsets or use an average across all offsets.
We compare the zero-offset to the mean after averaging over all subpixel offsets in Figure 8, finding that the average is within a few percent of the zero-shift value for most filters. However, the most under-sampled filters – namely F062, F087, and F106 – show that the average value is up to 40% smaller than the zero-shift value. For simulated observations using the bluest filters, our current predictions for position uncertainties may be pessimistic at the tens of percent level, but how that effect propagates into PM and parallax precision depends on the particular observing strategy.
After accounting for subpixel dithers, it is not surprising that smaller width PSFs could enable more precise position measurements compared to wider PSFs. Of course, this “narrower PSF gives more precision” trend is only true up to a certain point: for a delta function PSF, localizing a source is only possible to within approximately pixel. Using similar logic, under-sampled PSFs necessarily rely more on the information in their relatively low S/N wings when constraining subpixel positions. This likely explains why previous empirical calibrations with real HST , JWST , and Euclid observations have found that under-sampled PSFs do not localize stellar centroids as well as well-sampled ones at a given S/N (e.g., Bellini et al., 2011; Libralato et al., 2023, 2024a). Similarly, we expect that future real Roman data will reveal that the very under-sampled filters will measure less precise positions for the same flux SNR compared to better-sampled filters. While it is difficult to predict the magnitude of this effect, it further justifies our choice to use the more pessimistic zero-shift values instead of the subpixel-offset averaged ones.
The calculations presented in this work assume that there will not be significant uncertainties associated with aligning Roman images onto the Gaia reference frame, either through systematics or by simple propagation that leads to inflated localization uncertainties. That is, we assume the alignment uncertainty is a component of the floor uncertainty in our positions. Because of the large Roman field of view, this is likely a very reasonable choice: most pointings will contain enough well-measured Gaia stars (e.g., ) to align each image. However, there may be sparse regions – for example, in the HLWAS – where there are not many Gaia stars per Roman pointing. When properly propagated, this alignment uncertainty may increase the uncertainty of the Roman positions above what we have assumed.
@@ IV.3 Astrometry Calculations
When combining data from multiple observatories, systematic errors from all telescopes can and will manifest in the final combined astrometric solutions. While we have already discussed systematics in the Roman position measurements, previous work has demonstrated biases, correlations, and underestimated uncertainties in Gaia PMs and parallaxes (e.g., Vasiliev, 2019; Lindegren et al., 2021). We have not included Gaia -level systematics effects in this work because it is difficult to anticipate their strength or importance in future data releases. Ideally given enough time, the Gaia systematics may be fully characterized or removed entirely. One could include Gaia systematic effects by adding a floor astrometric covariance matrix to in Equation 3, although we note that this is not a feature of the current version of our simulation tool.